Write a Matlab code to generate Gold code and test its cross and auto correlation properties
Answer #1

# Make a shift register of length 10

# initilize with all 1's

G1 = [1 for i in xrange(10)]

print G1

fb3 = G1[3-1] # position 3

fb10 = G1[10-1]   # position 10

fb = (fb3 + fb10) % 2 # modulo 2 add

print fb

# shift everything one place to the right

for i in reversed(range(len(G1[1:]))):

G1[i+1] = G1[i]

# feedback

G1[0] = fb

print G1

fb = (G1[3-1] + G1[10-1]) % 2

for i in reversed(range(len(G1[1:]))):

G1[i+1] = G1[i]

G1[0] = fb

print G1

def shift(register, feedback, output):

"""GPS Shift Register

:param list feedback: which positions to use as feedback (1 indexed)

:param list output: which positions are output (1 indexed)

:returns output of shift register:

"""

# calculate output

out = [register[i-1] for i in output]

if len(out) > 1:

out = sum(out) % 2

else:

out = out[0]

# modulo 2 add feedback

fb = sum([register[i-1] for i in feedback]) % 2

# shift to the right

for i in reversed(range(len(register[1:]))):

register[i+1] = register[i]

# put feedback in position 1

register[0] = fb

return out

# example:

print shift(G1, [3,10], [10])

print G1